OpenStudy (anonymous):

Graph the line that goes through point (–5, 5) with slope 647-03-05-00-00_files/i0150000.jpg.

OpenStudy (luigi0210):

try attaching the document again?

OpenStudy (anonymous):

(-5,5)with slope 1/5 Y=mX+b 5=1/5*-5+b 5=-5/5+b 5=-1+b b=5+1 b=6 Y=X/5+6 from -10 to 10, x/5 +6).